# Orbital Optical Raman Lattice

### v2

@@ Abstract
The spin and orbital are two basic degrees of freedom, which play significant roles in exploring exotic quantum phases in optical lattices with synthetic spin-orbit coupling (SOC) and high orbital bands, respectively. Here, we propose an orbital optical Raman lattice to study exotic high-orbital Bose condensates with Raman-induced SOC in a square lattice. We find that both the SOC and $p$-orbital interactions influence the condensed state of bosons. Their interplay results in two novel high-orbital many-body quantum phases: the uniform angular momentum superfluid phase, which features a global topological chiral orbital current, and the two-dimensional spin-orbital supersolid phase, which is characterized by the spin and orbital angular momentum density wave patterns. The former has topological Bogoliubov excitations characterized by a uniform Chern number, while the spin-orbital supersolid phase hosts topological excitations with opposite Chern numbers which respectively protect the chiral and antichiral edge modes in the neighboring supersolid clusters. Our findings may inspire a new avenue for exploring exotic SOC and high-orbital physics in optical lattices, and shall also advance the experimental progress of novel supersolids in higher dimensions.
@@ Introduction.
Ultracold atomic gases provide ideal platforms for quantum simulation due to their pristine nature and full controllability [1, 2]. Among them, investigations into synthetic gauge fields [3, 4, 5, 6, 7, 8, 9, 10, 11] and spin-orbit couplings (SOCs) [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] have attracted widespread interests. Over the past decade, experimental realizations of novel SOCs in one-dimension [14, 17, 18, 24, 25, 26], two-dimension [23, 29, 30, 31, 32], and three-dimension [33] have enabled the simulation of various exotic topological models with cold atoms, such as topological semimetals [33], quantum anomalous Hall insulator [34], and non-Hermitian topological phases [35]. Additionally, extending synthetic SOC to strongly correlated regime facilitates the realization of non-Abelian dynamical gauge fields [36]. Besides simulating topological quantum phases [37, 38, 39, 40, 41, 42, 43], the quest for supersolids characterized by diagonal and off-diagonal long-range order in ultracold atoms has never ceased [44, 45, 46]. Currently, the supersolid phase has been observed in long-range interacting dipolar quantum gases [47, 48, 49, 50, 51, 52, 53] and optical cavities [54, 55].
@@
High-orbital (e.g. $p$ and $d$) systems in optical lattices also received considerable attention due to their rich degrees of freedom, which can give rise to exotic orbital physics [56, 57, 58, 59, 60, 61, 62, 63]. With the orbital degree of freedom, one can not only simulate the behavior of electrons in realistic materials but also, more crucially, uncover new concepts and phenomena that have no prior analog in electronic systems [63], such as the high-orbital Bose-Einstein condensates (BECs) with novel orbital ordering [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. To date, significant progresses have been made in the study of $p$-orbital BEC in optical lattices. For instance, the $p$-orbital BEC in a square lattice manifests as a superfluid phase with a staggered angular momentum order that breaks time-reversal symmetry [66]. In a triangular lattice, it results in staggered loop current orders [67, 79], and in a hexagonal lattice, it exhibits Potts-nematic superfluid order [76] or atomic chiral superfluidity with topological Bogoliubov excitations [77]. Such phases reveal the remarkable richness of higher orbital coherence whether or not breaking the time-reversal symmetry.
@@
Figure 1: (a) Generic setup for orbital optical Raman lattice. The bias magnetic field $\bm{B}$ is parallel to the x direction. A pair of standing waves, $\bm{E}_{xy}$ and $\bm{E}_{yz}$, produce a square lattice. Moreover, the Raman field can be achieved by applying another plane wave $\bm{E}_{zx}$. (b) Optical-dipole transition diagram for Raman coupling in cold bosons (${}^{87}\mathrm{Rb}$) coupled to two pairs of laser beams ($\bm{E}_{xy}$, $\bm{E}_{zx}$) and ($\bm{E}_{yz}$, $\bm{E}_{zx}$). Here the hyperfine states $|1,-1\rangle$ and $|1,0\rangle$ of ${}^{87}\mathrm{Rb}$ are used to emulate the ground states $|g_{\uparrow,\downarrow}\rangle$. (c) The sketch for the nearest-neighbor spin-conserved hopping for $s$ and $p_{x,y}$ orbitals and the spin-flipped hopping between $s$ and $p_{x,y}$. Here, Only the $p_{x}$ orbital is displayed, since the hopping terms for $p_{x}$ and $p_{y}$ are related by $C_{4}$ symmetry. (d) Single-particle spectrum. $\bm{k}_{0}^{1}$, $\bm{k}_{0}^{2}$, $\bm{k}_{0}^{3}$, $\bm{k}_{0}^{4}$ are the lowest energy points connected by $C_{4}$ symmetry. Thereinafter, $V_{0}=5.0E_{r}$, $M_{0}=1.0E_{r}$ (tight-binding parameters: $t_{s}=0.0658E_{r}$, $t_{p}^{\parallel}=0.4228E_{r}$, $t_{p}^{\perp}=0.0658E_{r}$, $t_{so}=0.1059E_{r}$, $\mu_{s}=5.1909E_{r}$, and $\mu_{p}=5.9049E_{r}$), $m_{z}=0.0173E_{r}$. Besides, $\bm{k}_{0}^{1}=(0.154,0.154)\pi$. Note that $E_{r}$ is the recoil energy.
@@
In this letter, we propose an orbital optical Raman lattice that combines the both topics of synthetic SOC [23, 21, 28] and $p$-orbital condensates [66] in a square lattice, and predict two novel quantum many-body phases with nontrivial topology: the uniform angular momentum superfluid (UAMSF) and the 2D spin-orbital supersolid (SOSS), as characterized by different spin and orbital angular momentum orders. The condensed state of bosons is shown to be governed by the nontrivial interplay effects of the on-site $p$-orbital interactions and Raman-induced SOC. For the $p$-orbital interaction dominating over the Raman-induced SOC, the ground state exhibits a uniform angular momentum order, dubbed the UAMSF, which hosts uniform chiral topological excitations. In contrast, when the $p$-orbital interaction is comparable to SOC, the staggered spin and angular momentum cluster patterns emerge, leading to the 2D topological SOSS which host chiral and antichiral topological edge excitations in the neighboring supersolid clusters. Our work offers a new perspective for exploring exotic high-orbital physics with nontrivial SOC effects and shall advance the realization of novel supersolid phases without long-range interactions in ultracold atom systems.
@@
We start with the Hamiltonian for ultracold bosons trapped in an orbital optical Raman lattice, given by $H_{0}=\frac{p_{x}^{2}}{2m}+\frac{p_{y}^{2}}{2m}+V(x,y)+M(x,y)\sigma_{x}+m_{z}\sigma_{z}$. Here, $V(x,y)=-V_{0}[\cos^{2}(\frac{\pi}{a}x)+\cos^{2}(\frac{\pi}{a}y)]$ and $M(x,y)=-M_{0}[\cos(\frac{\pi}{a}x)+\cos(\frac{\pi}{a}y)]$ represent the normal lattice potential and Raman field, respectively. The Pauli matrices $\bm{\sigma}$ act on the subspace spanned by $|s\rangle\otimes|g_{\uparrow}\rangle$ and $|p\rangle\otimes|g_{\downarrow}\rangle$, where the spin $|g_{\uparrow,\downarrow}\rangle$ is defined from hyperfine ground states, $|s\rangle$ and $|p\rangle$ represent local orbital states. The model Hamiltonian $H_{0}$ can be realized with high feasibility, through the three-laser configuration shown in Fig. 1(a). Two lasers form standing waves, $\bm{E}_{xy}=\hat{y}E_{0}\cos(k_{0}x)$ and $\bm{E}_{yz}=\hat{z}E_{0}\cos(k_{0}y)$, linearly polarized along $y$ and $z$ axes, respectively, and propagating in $x$-$y$ plane. The third laser forms a plane wave, $\bm{E}_{zx}=\hat{x}E_{z}e^{-\mathrm{i}k_{z}z}$ propagating along $z$ direction and is $x$-polarized. The standing waves $\bm{E}_{xy}$ and $\bm{E}_{yz}$ coupling to excited states with a red detuning $\Delta$ contribute to the diagonal square optical lattice potentials $V(x,y)=\hbar(|\bm{E}_{xy}|^{2}/\Delta+|\bm{E}_{yz}|^{2}/\Delta)$. Additionally, $\bm{E}_{xy}$, $\bm{E}_{yz}$, and $\bm{E}_{zx}$ induce the Raman field components $M(x)$ and $M(y)$ via two two-photon transitions as illustrated in a $\bm{\varLambda}$-type configuration [Fig.1(b)]. In experiment one can easily set that the Raman field only couples $|s\rangle\otimes|g_{\uparrow}\rangle$ and $|p\rangle\otimes|g_{\downarrow}\rangle$ by putting such two states be nearly resonant for the two-photon transitions except for a tunable two-photon detuning $\delta$, which defines an effective Zeeman splitting $m_{z}=\hbar\delta/2$, while all other orbital states are far detuned [80].
@@
With the above implementation scheme, bosons can occupy the spin-orbital locking states ϕssubscriptitalic-$\phi_{s_{\uparrow}}$, $\phi_{p_{x\downarrow}}$, and $\phi_{p_{y\downarrow}}$. As shown in Fig.1(c), the lattice potential $V(x,y)$ governs spin-conserved nearest-neighbour hopping ($t_{s}$ and $t_{p}$), while the Raman field $M(x,y)$ contributes to spin-flipped nearest-neighbour hopping ($t_{so}$). Since the Raman field $M(x,y)$ has twice the period of the optical lattice potential $V(x,y)$ and is symmetric with respect to each lattice site center of $V(x,y)$, the spin-flipped hopping is staggered in the $\nu$ ($\nu=x,y$) direction with $t_{so}^{(\bm{i},\bm{i}\pm\bm{e}_{v})}=\pm(-1)^{i_{v}}t_{\mathrm{SO}}$ [80]. The staggered factor $(-1)^{i_{\nu}}$ represents a nontrivial momentum transfer between spin-up and spin-down states in the Raman coupling, and can be absorbed by the transformation $b_{\bm{i},p_{\nu},\downarrow}\rightarrow-(-1)^{i_{\nu}}b_{\bm{i},p_{\nu},\downarrow}$, yielding the tight-binding Hamiltonian
@@
where $b_{\bm{i},l,\sigma}^{\dagger}$ ($b_{\bm{i},l,\sigma}$) ($l=s,p_{x},p_{y}$) denotes the creation (annihilation) operator of boson. $\bar{t}_{p}^{\mu\nu}=t_{p}^{\parallel}\delta_{\mu\nu}+t_{p}^{\perp}(1-\delta_{\mu\nu})$ ($\mu$, $\nu$=$x$, $y$), $n_{\bm{i},l,\sigma}=b_{\bm{i},l,\sigma}^{\dagger}b_{\bm{i},l,\sigma}$, and $\mu_{s,p}$ denote the on-site energies. A key feature is that the momentum transfer by Raman field forces the band minima of $\epsilon_{p_{x}\downarrow}(\bm{k})$ ($\epsilon_{p_{y}\downarrow}(\bm{k})$) to move from $\bm{Q}_{x}=(\pi,0)$ ($\bm{Q}_{y}=(0,\pi)$) to $\bm{\Gamma}$ [80]. Further, the remaining 2D SOC effect can result in a single-particle spectrum with four $C_{4}$-connected minima at $\bm{k}_{0}^{1}$, $\bm{k}_{0}^{2}$, $\bm{k}_{0}^{3}$, and $\bm{k}_{0}^{4}$ [Fig.1(d)], leading to the major novel physics predicted in the present orbital optical Raman lattice. A quadratic band touch point protected by time-reversal symmetry exists at $\bm{\Gamma}$ [81].
Figure 2: (a) Schematic of the variational calculation. Here, $\bm{k}_{0}^{n}$, $\bm{k}_{c}^{n}$, and $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) represent the energy minima of the single-particle spectrum, the variational condensation momenta under weak interaction, and the real condensation momenta, respectively. (b) Ground state phase diagram. Here, $m_{z}$ and $n_{0}$ denote the effective Zeeman splitting and boson condensation density, respectively. There are three types of ground states: $C_{4}$-symmetry-broken superfluid phase ($C_{4}$-SBSF), two-dimensional spin-orbital supersolid phase (2D SOSS), and uniform angular momentum superfluid phase (UAMSF). Parameters: $V_{0}=5.0E_{r}$, $M_{0}=1.0E_{r}$, and $g=0.01E_{r}$ ($U_{s}=0.1044E_{r}$, $U_{sp}=0.0428E_{r}$, $U_{p}=0.0618E_{r}$, and $\widetilde{U}_{p}=0.0175E_{r}$).
The total Hamiltonian includes also the contact interaction $V(\bm{r}-\bm{r}^{\prime})=g_{\sigma\sigma^{\prime}}\delta(\bm{r}-\bm{r}^{\prime})$ ($\sigma,\sigma^{\prime}=\uparrow,\downarrow$), whose form in the lattice model reads [80]
@@
Usually, the non-interacting BEC will take place at the single-particle band minimum points $\bm{k}_{0}^{n}$ ($n=1\sim 4$). However, since these lowest energy points of Hamiltonian $\hat{H}_{\mathrm{0}}$ do not generally correspond to the energy minima of the total Hamiltonian $\hat{H}_{\mathrm{0}}+\hat{H}_{\mathrm{int}}$, the interaction changes the populations in the $s_{\uparrow}$, $p_{x\downarrow}$, and $p_{y\downarrow}$ orbitals, which shifts the condensation momenta around $\bm{k}_{0}^{n}$ ($n=1\sim 4$). Under weak interaction, the condensed ground state ansatz can be given by $|g\rangle\sim e^{\sqrt{N_{0}}b^{\dagger}}|\mathrm{vac}\rangle$, where $b^{\dagger}$, expressed as $b^{\dagger}=\sum_{n=1}^{4}\sum_{l_{\sigma}}\gamma_{\bm{k}_{c}^{n}}\beta_{\bm{k}_{c}^{n},l_{\sigma}}b_{\bm{k}_{c}^{n},l_{\sigma}}^{\dagger}$ ($l_{\sigma}=s_{\uparrow},p_{x\downarrow},p_{y\downarrow}$), represents the quasi-particle creation operator and $N_{0}$ is the particle number of BEC. Here, $C_{4}$-related points $\bm{k}_{c}^{n}$ [Fig.2(a)] denote the variational condensation momentum, with the variational parameters $\gamma_{\bm{k}_{c}^{n}}$ and $\beta_{\bm{k}_{c}^{n}}$ satisfying $\sum_{n}|\gamma_{\bm{k}_{c}^{n}}|^{2}=1$ and $\sum_{l_{\sigma}}|\beta_{\bm{k}_{c}^{n},l_{\sigma}}|^{2}=1$, respectively. Note that $|\gamma_{\bm{k}_{c}^{n}}|^{2}$ represents the condensation distribution probability of bosons at $\bm{k}_{c}^{n}$, and $|\beta_{\bm{k}_{c}^{n},l_{\sigma}}|^{2}$ denotes the orbital population probability. To identify the true condensation momenta $\tilde{\bm{k}}_{c}^{n}$ and the condensation parameters ($\gamma_{\tilde{\bm{k}}_{c}^{n}}$, $\beta_{\tilde{\bm{k}}_{c}^{n},l_{\sigma}}$), we search for the minimum of the energy density functional $\langle g|\hat{\mathcal{H}}_{0}+\hat{\mathcal{H}}_{\mathrm{int}}|g\rangle/N$ by simulated annealing algorithm [82], where $N$ is the number of lattice sites. As depicted in Fig.2(a), $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) lie on the diagonal lines and move toward $\bm{\Gamma}$ (or coincide with it), and they are also connected by $C_{4}$ symmetry.

### v3

@@ Abstract
Spin and orbital are two basic degrees of freedom that play significant roles in exploring exotic quantum phases in optical lattices with synthetic spin-orbit coupling (SOC) and high orbital bands, respectively. Here, we combine these two crucial ingredients for the first time by proposing a completely new orbital optical Raman lattice scheme to explore exotic high-orbital Bose condensates with Raman-induced SOC in a square lattice. We find that both the SOC and $p$-orbital interactions influence the condensed state of bosons. Their interplay results in two novel high-orbital many-body quantum phases: the uniform angular momentum superfluid phase, which exhibits a global topological chiral orbital current characterized by a uniform Chern number, and the two-dimensional topological spin-orbital supersolid phase, which is characterized by the spin and orbital angular momentum density wave patterns and topological excitations with opposite Chern numbers, respectively protecting the chiral and antichiral edge modes in the neighboring supersolid clusters. Our scheme may open a new avenue for exploring exotic SOC and high-orbital physics in optical lattices, and is expected to advance the experimental realization of novel supersolids in higher dimensions.
@@ Introduction.
Ultracold atomic gases provide ideal platforms for quantum simulation due to their pristine nature and full controllability Bloch et al. (2008); Gross and Bloch (2017). Among them, investigations into synthetic gauge fields Jaksch and Zoller (2003); Juzeliūnas and Öhberg (2004); Osterloh et al. (2005); Ruseckas et al. (2005); Liu et al. (2006); Dalibard et al. (2011); Goldman et al. (2014); Chen et al. (2018); Aidelsburger et al. (2022) and spin-orbit couplings (SOCs) Liu et al. (2009); Juzeliūnas et al. (2010); Lin et al. (2011); Campbell et al. (2011); Sau et al. (2011); Wang et al. (2012); Cheuk et al. (2012); Anderson et al. (2012); Galitski and Spielman (2013); Liu et al. (2014); Zhai (2015); Wu et al. (2016); Burdick et al. (2016); Song et al. (2016); Chen et al. (2016); Kolkowitz et al. (2017); Zhang and Liu (2018); Wang et al. (2018) have attracted widespread interests. Over the past decade, experimental realizations of novel SOCs in one-dimension Lin et al. (2011); Wang et al. (2012); Cheuk et al. (2012); Burdick et al. (2016); Song et al. (2016); Kolkowitz et al. (2017), two-dimension Wu et al. (2016); Huang et al. (2016); Meng et al. (2016); Sun et al. (2018a, b), and three-dimension Wang et al. (2021a) have enabled the simulation of various exotic topological models with cold atoms, such as topological semimetals Wang et al. (2021a), quantum anomalous Hall insulator Liang et al. (2023), and non-Hermitian topological phases Zhao et al. (2023). Additionally, extending synthetic SOC to strongly correlated regime facilitates the realization of non-Abelian dynamical gauge fields Zhou et al. (2023). Besides simulating topological quantum phases Goldman et al. (2010); Béri and Cooper (2011); Liu et al. (2013); Jotzu et al. (2014); Wang et al. (2014); Zhang and Zhou (2017); Song et al. (2018); Yi et al. (2019), the quest for supersolids characterized by diagonal and off-diagonal long-range order in ultracold atoms has never ceased Boninsegni and Prokof’ev (2012); Ritsch et al. (2013); Recati and Stringari (2023). Currently, the supersolid phase has been observed in long-range interacting dipolar quantum gases Li et al. (2017); Böttcher et al. (2019); Tanzi et al. (2019); Chomaz et al. (2019); Norcia et al. (2021); Bland et al. (2022); Casotti et al. (2024) and optical cavities Léonard et al. (2017a, b).
@@
High-orbital (e.g. $p$ and $d$) systems in optical lattices also received considerable attention due to their rich degrees of freedom, which can give rise to exotic orbital physics Zhao and Liu (2008); Liu et al. (2010); Cai et al. (2011); Lewenstein and Liu (2011); Li et al. (2013); Sowiński et al. (2013); Dutta et al. (2015); Li and Liu (2016); Kock et al. (2016). With the orbital degree of freedom, one can not only simulate the behavior of electrons in realistic materials but also, more crucially, uncover new concepts and phenomena that have no prior analog in electronic systems Li and Liu (2016), such as the high-orbital Bose-Einstein condensates (BECs) with novel orbital ordering Isacsson and Girvin (2005); Kuklov (2006); Liu and Wu (2006); Wu et al. (2006); Larson et al. (2009); Wirth et al. (2011); Li et al. (2012); Xu et al. (2016); Di Liberto et al. (2016); Wang and Liu (2017); Li et al. (2018); Niu et al. (2018); Pan et al. (2020); Jin et al. (2021); Wang et al. (2021b); Huang et al. (2022); Wang et al. (2023). To date, significant progress have been made in the study of $p$-orbital BEC in optical lattices. For instance, the $p$-orbital BEC in a square lattice manifests as a superfluid phase with a staggered angular momentum order that breaks time-reversal symmetry Liu and Wu (2006). In a triangular lattice, it results in staggered loop current orders Wu et al. (2006); Wang et al. (2023), and in a $sp^{2}$-orbital hexagonal lattice, it exhibits Potts-nematic superfluid order Jin et al. (2021) or atomic chiral superfluidity with topological excitations Wang et al. (2021b). Such phases reveal the remarkable richness of higher orbital coherence whether or not breaking the time-reversal symmetry.
@@
Figure 1: (a) Generic setup for orbital optical Raman lattice. The bias magnetic field $\bm{B}$ is parallel to the x direction. A pair of standing waves, $\bm{E}_{xy}$ and $\bm{E}_{yz}$, produce a square lattice. Moreover, the Raman field can be achieved by applying another plane wave $\bm{E}_{zx}$. (b) Optical-dipole transition diagram for Raman coupling in cold bosons (${}^{87}\mathrm{Rb}$) coupled to two pairs of laser beams ($\bm{E}_{xy}$, $\bm{E}_{zx}$) and ($\bm{E}_{yz}$, $\bm{E}_{zx}$). Here the hyperfine states |1,1|1,-1$\rangle$| 1 , - 1 ⟩ and |1,0|1,0$\rangle$| 1 , 0 ⟩ of ${}^{87}\mathrm{Rb}$ are used to emulate the ground states $|g_{\uparrow,\downarrow}\rangle$. (c) The sketch for the nearest-neighbor spin-conserved hopping for $s$ and $p_{x,y}$ orbitals and the spin-flipped hopping between $s$ and $p_{x,y}$. Here, only the $p_{x}$ orbital is displayed, since the hopping terms for $p_{x}$ and $p_{y}$ are related by $C_{4}$ symmetry. (d) Single-particle spectrum. The $C_{4}$-related points $\bm{k}_{0}^{1}$, $\bm{k}_{0}^{2}$, $\bm{k}_{0}^{3}$, $\bm{k}_{0}^{4}$ are the lowest energy points that lie on the diagonal lines in the Brillouin zone. Thereinafter, $V_{0}=5.0E_{r}$, $M_{0}=1.0E_{r}$ (tight-binding parameters: $t_{s}=0.0658E_{r}$, $t_{p}^{\parallel}=0.4228E_{r}$, $t_{p}^{\perp}=0.0658E_{r}$, $t_{so}=0.1059E_{r}$, $\mu_{s}=5.1909E_{r}$, and $\mu_{p}=5.9049E_{r}$), $m_{z}=0.0173E_{r}$. Besides, $\bm{k}_{0}^{1}=(0.154,0.154)\pi$. $E_{r}$ is the recoil energy.
@@
In this letter, we propose an orbital optical Raman lattice scheme that integrates the synthetic SOC Wu et al. (2016); Liu et al. (2014); Wang et al. (2018) and $p$-orbital condensates Isacsson and Girvin (2005); Kuklov (2006); Liu and Wu (2006), predicting two exotic quantum many-body phases with nontrivial interaction-induced topology in the square lattice: the uniform angular momentum superfluid (UAMSF) and the 2D topological spin-orbital supersolid (SOSS), both of which go beyond the scope of previous studies on synthetic SOC and high-orbital physics. We show that these condensed states of bosons are governed by the nontrivial interplay effects between on-site $p$-orbital interactions and Raman-induced SOC, surpassing the limitations of pure orbital lattice gas models without synthetic SOC. The UAMSF exhibits a uniform angular momentum order, supporting uniform chiral topological excitations. In contrast, the 2D topological SOSS shows staggered spin and angular momentum cluster patterns, leading to the chiral and antichiral topological edge excitations in the neighboring supersolid clusters. Our work offers a new perspective for the combined study of nontrivial synthetic SOC and high-orbital physics, which were previously considered separate fields in optical lattice research, and will promote the realization of novel supersolids without long-range interactions in ultracold atom systems.
@@
We start with the Hamiltonian for ultracold bosons trapped in an orbital optical Raman lattice, given by $H_{0}=\frac{p_{x}^{2}}{2m}+\frac{p_{y}^{2}}{2m}+V(x,y)+M(x,y)\sigma_{x}+m_{z}\sigma_{z}$. Here, $V(x,y)=-V_{0}[\cos^{2}(\frac{\pi}{a}x)+\cos^{2}(\frac{\pi}{a}y)]$ and $M(x,y)=-M_{0}[\cos(\frac{\pi}{a}x)+\cos(\frac{\pi}{a}y)]$ represent the normal lattice potential and Raman field, respectively. The Pauli matrices $\bm{\sigma}$ act on the subspace spanned by $|s\rangle\otimes|g_{\uparrow}\rangle$ and $|p\rangle\otimes|g_{\downarrow}\rangle$, where the spin $|g_{\uparrow,\downarrow}\rangle$ is defined from hyperfine ground states, $|s\rangle$ and $|p\rangle$ represent local orbital states. The model Hamiltonian $H_{0}$ can be realized with high feasibility, through the three-laser configuration shown in Fig. 1(a). Two lasers form standing waves, $\bm{E}_{xy}=\hat{y}E_{0}\cos(k_{0}x)$ and $\bm{E}_{yz}=\hat{z}E_{0}\cos(k_{0}y)$, linearly polarized along $y$ and $z$ axes, respectively, and propagating in $x$-$y$ plane. The third laser forms a plane wave, $\bm{E}_{zx}=\hat{x}E_{z}e^{-\mathrm{i}k_{z}z}$ propagating along $z$ direction and is $x$-polarized. The standing waves $\bm{E}_{xy}$ and $\bm{E}_{yz}$ coupling to excited states with a red detuning $\Delta$ contribute to the diagonal square optical lattice potentials $V(x,y)=\hbar(|\bm{E}_{xy}|^{2}/\Delta+|\bm{E}_{yz}|^{2}/\Delta)$. Additionally, $\bm{E}_{xy}$, $\bm{E}_{yz}$, and $\bm{E}_{zx}$ induce the Raman field components $M(x)$ and $M(y)$ via two two-photon transitions as illustrated in a $\bm{\varLambda}$-type configuration [Fig.1(b)]. In experiment one can easily set that the Raman field only couples $|s\rangle\otimes|g_{\uparrow}\rangle$ and $|p\rangle\otimes|g_{\downarrow}\rangle$ by putting such two states be nearly resonant for the two-photon transitions except for a tunable two-photon detuning $\delta$, which defines an effective Zeeman splitting $m_{z}=\hbar\delta/2$, while all other orbital states are far detuned SM .
@@
With the above implementation scheme, bosons can occupy the spin-orbital locking states $\phi_{s_{\uparrow}}$, $\phi_{p_{x\downarrow}}$, and $\phi_{p_{y\downarrow}}$. As shown in Fig.1(c), the lattice potential $V(x,y)$ governs spin-conserved nearest-neighbour hopping ($t_{s}$ and $t_{p}$), while the Raman field $M(x,y)$ contributes to spin-flipped nearest-neighbour hopping ($t_{so}$). Since the Raman field $M(x,y)$ has twice the period of the optical lattice potential $V(x,y)$ and is symmetric with respect to each lattice site center of $V(x,y)$, the spin-flipped hopping is staggered in the $\nu$ ($\nu=x,y$) direction with $t_{so}^{(\bm{i},\bm{i}\pm\bm{e}_{v})}=\pm(-1)^{i_{v}}t_{\mathrm{SO}}$ SM . The staggered factor $(-1)^{i_{\nu}}$ represents a nontrivial momentum transfer between spin-up and spin-down states in the Raman coupling, and can be absorbed by the transformation $b_{\bm{i},p_{\nu},\downarrow}\rightarrow-(-1)^{i_{\nu}}b_{\bm{i},p_{\nu},\downarrow}$, yielding the tight-binding Hamiltonian
@@
where $b_{\bm{i},l,\sigma}^{\dagger}$ ($b_{\bm{i},l,\sigma}$) ($l=s,p_{x},p_{y}$) denotes the creation (annihilation) operator of boson. $\bar{t}_{p}^{\mu\nu}=t_{p}^{\parallel}\delta_{\mu\nu}+t_{p}^{\perp}(1-\delta_{\mu\nu})$ ($\mu$, $\nu$=$x$, $y$), $n_{\bm{i},l,\sigma}=b_{\bm{i},l,\sigma}^{\dagger}b_{\bm{i},l,\sigma}$, and $\mu_{s,p}$ denote the on-site energies. A key feature is that the momentum transfer by Raman field forces the band minima of $\epsilon_{p_{x}\downarrow}(\bm{k})$ ($\epsilon_{p_{y}\downarrow}(\bm{k})$) to move from $\bm{Q}_{x}=(\pi,0)$ ($\bm{Q}_{y}=(0,\pi)$) to $\bm{\Gamma}$ SM . Further, the remaining 2D SOC effect can result in a single-particle spectrum with four $C_{4}$-connected minima at $\bm{k}_{0}^{1}$, $\bm{k}_{0}^{2}$, $\bm{k}_{0}^{3}$, and $\bm{k}_{0}^{4}$ [Fig.1(d)], leading to the major novel physics predicted in the present orbital optical Raman lattice. A quadratic band touch point protected by time-reversal symmetry exists at $\bm{\Gamma}$ Sun et al. (2012).
Figure 2: (a) Schematic of the variational calculation. Here, $C_{4}$-connected points $\bm{k}_{0}^{n}$, $\bm{k}_{c}^{n}$, and $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) represent the energy minima of the single-particle spectrum, the variational momenta within the areas surrounded by red circles, and the condensation momenta along the blue diagonal lines, respectively. (b) Ground state phase diagram. $m_{z}$ and $n_{0}$ denote the effective Zeeman splitting and boson condensation density, respectively. There are three types of ground states: $C_{4}$-symmetry-broken superfluid phase ($C_{4}$-SBSF), two-dimensional spin-orbital supersolid phase (2D SOSS), and uniform angular momentum superfluid phase (UAMSF). Parameters: $V_{0}=5.0E_{r}$, $M_{0}=1.0E_{r}$, and $g=0.01E_{r}$ ($U_{s}=0.1044E_{r}$, $U_{sp}=0.0428E_{r}$, $U_{p}=0.0618E_{r}$, and $\widetilde{U}_{p}=0.0175E_{r}$).
The total Hamiltonian includes also the contact interaction $V(\bm{r}-\bm{r}^{\prime})=g_{\sigma\sigma^{\prime}}\delta(\bm{r}-\bm{r}^{\prime})$ ($\sigma,\sigma^{\prime}=\uparrow,\downarrow$), whose form in the lattice model reads SM
@@
Usually, the non-interacting BEC will take place at the single-particle band minimum points $\bm{k}_{0}^{n}$ ($n=1\sim 4$). However, since these lowest energy points of Hamiltonian $\hat{H}_{\mathrm{0}}$ do not generally correspond to the energy minima of the total Hamiltonian $\hat{H}_{\mathrm{0}}+\hat{H}_{\mathrm{int}}$, the interaction changes the populations in the $s_{\uparrow}$, $p_{x\downarrow}$, and $p_{y\downarrow}$ orbitals, which shifts the condensation momenta around $\bm{k}_{0}^{n}$ ($n=1\sim 4$). Under weak interaction, the condensed ground state ansatz can be given by $|g\rangle\sim e^{\sqrt{N_{0}}b^{\dagger}}|\mathrm{vac}\rangle$, where $b^{\dagger}$, expressed as $b^{\dagger}=\sum_{n=1}^{4}\sum_{l_{\sigma}}\gamma_{\bm{k}_{c}^{n}}\beta_{\bm{k}_{c}^{n},l_{\sigma}}b_{\bm{k}_{c}^{n},l_{\sigma}}^{\dagger}$ ($l_{\sigma}=s_{\uparrow},p_{x\downarrow},p_{y\downarrow}$), represents the quasi-particle creation operator and $N_{0}$ is the particle number of BEC. Here, $C_{4}$-related points $\bm{k}_{c}^{n}$ [Fig.2(a)] denote the variational momenta, with the variational parameters $\gamma_{\bm{k}_{c}^{n}}$ and $\beta_{\bm{k}_{c}^{n}}$ satisfying $\sum_{n}|\gamma_{\bm{k}_{c}^{n}}|^{2}=1$ and $\sum_{l_{\sigma}}|\beta_{\bm{k}_{c}^{n},l_{\sigma}}|^{2}=1$, respectively. Note that $|\gamma_{\bm{k}_{c}^{n}}|^{2}$ represents the condensation distribution probability of bosons at $\bm{k}_{c}^{n}$, and $|\beta_{\bm{k}_{c}^{n},l_{\sigma}}|^{2}$ denotes the orbital population probability. To identify the condensation momenta $\tilde{\bm{k}}_{c}^{n}$ and the condensation parameters ($\gamma_{\tilde{\bm{k}}_{c}^{n}}$, $\beta_{\tilde{\bm{k}}_{c}^{n},l_{\sigma}}$), we search for the minimum of the energy density functional $\langle g|\hat{\mathcal{H}}_{0}+\hat{\mathcal{H}}_{\mathrm{int}}|g\rangle/N$ by simulated annealing algorithm, with $N$ the number of lattice sites [see Supplementary for details SM ]. As depicted in Fig.2(a), $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) lie on the blue diagonal lines and move toward $\bm{\Gamma}$ (or coincide with it), and are connected by $C_{4}$ symmetry.